ABC and CVP Analysis: Multiple Products Good Scent, Inc., produces two colognes: Rose and Violet. Of the two, Rose is more popular. Data concerning the two products follow: Rose Violet Expected sales (in cases) 54,500 10,900 Selling price per case $98 $83 Direct labor hours 33,550 5,800 Machine hours 10,050 3,100 Receiving orders 53 24 Packing orders 97 48 Material cost per case $50 $43 Direct labor cost per case $11 $9 The company uses a conventional costing system and assigns overhead costs to products using direct labor hours. Annual overhead costs follow. They are classified as fixed or variable with respect to direct labor hours. Fixed Variable Direct labor benefits $ — $181,010 Machine costs 198,000* 251,840 Receiving department 244,500 — Packing department 106,500 — Total costs $549,000 $432,850 * All depreciation Required: 1. Using the conventional approach, compute the number of cases of Rose and the number of cases of Violet that must be sold for the company to break even. In your computations, round variable unit cost to the nearest cent and round the number of break-even packages to the nearest whole number. Break-even cases of Rose Break-even cases of Violet 2. Using an activity-based approach, compute the number of cases of each product that must be sold for the company to break even. In your computations, round all computed amounts to the nearest cent and round the number of break-even packages to the nearest whole number. Break-even cases of Rose Break-even cases of Violet
Answer #1

Solution:

Solution 1) Calculation of Break Even Point in Units using Traditional Costing

Assignment of Overheads Costs using Direct Labour Hours

Total Direct Labour Hours = Direct Labour Hours of Rose + Direct Labour Hours of Voilet

                                                = 33550 + 5800 = 39350 Hours.

Direct Labour Benefits Cost assigned to Rose = (33550 / 39350) x $181,010 = $154,330

Direct Labour Benefits Cost assigned to Voilet = (5800 / 39350) x $181,010 = $26,680

Fixed Machine Cost assigned to Rose = (33550 / 39350) x $198000 = $168,815.76

Fixed Machine Cost assigned to Voilet = (5800 / 39350) x $198000 = $29184.24

Variable Machine Cost assigned to Rose = (33550 / 39350) x $251840 = $214720

Variable Machine Cost assigned to Voilet = (5800 / 39350) x $251840 = $37120

Receiving department Cost assigned to Rose = (33550 / 39350) x $244500 = $208461.88

Receiving department Cost assigned to Voilet = (5800 / 39350) x $244500 = $36038.12

Packing department Cost assigned to Rose = (33550 / 39350) x $244500 = $90802.41

Packing department Cost assigned to Voilet = (5800 / 39350) x $244500 = $15697.59

Income Statement as per Traditional Costing

Rose

Voilet

Total $ (54,500 Units)

Per Unit $

Total $ (10,900 Units)

Per Unit $

Sales

                    53,41,000.00

                                            98.00

                         9,04,700.00

                 83.00

Less: Variable Cost

Material

                    27,25,000.00

                                             50.00

                         4,68,700.00

                 43.00

Labour

                      5,99,500.00

                                             11.00

                             98,100.00

                   9.00

Direct Labour Benefits

                      1,54,330.00

                                               2.83

                             26,680.00

                   2.45

Machine Costs

2,14,720.00

                                               3.94

37,120.00

                   3.41

Total Variable Cost

                    36,93,550.00

                                             67.77

                         6,22,664.24

                 57.85

Contribution = Sales - Total Variable Cost

                    16,47,450.00

                                             30.23

                       16,47,450.00

                 25.15

Less: Fixed Cost (Overheads)

Machine Costs

                      1,68,815.76

                                               3.10

                             29,184.24

                   2.68

Receiving department

                      2,08,461.88

                                               3.82

                             36,038.12

                   3.31

Packing department

                          90,802.41

                                               1.67

                             15,697.59

                   1.44

Total Fixed Cost

                      4,68,080.05

                                               8.59

                             80,919.95

                   7.42

Profit = Contribution - Total Fixed Cost

                    11,79,369.95

                                             21.64

                       11,79,369.95

                 17.72

Calculation of Break Even Point (in Units) as per Traditional Costing

= Total Fixed Cost / Contribution per unit

For Rose =$468080.05 / $30.23 = 15,484.76 packages

For Voilet =$80,919.95 / $25.15 = 3,217.90 packages

Solution 2) Calculation of Break Even Point in Units using Activity Based Costing

Calculation of Rate per Cost Driver as per Activity Based Costing

Overheads

Cost $

Cost Driver

Number of Cost Drivers

Rate Per Cost Driver = Cost $ / Number of Cost Drivers

Variable Direct labour benefits $

                      1,81,010.00

Number of Direct Labour Hours

            39,350

                                           4.60

Fixed Machine Costs

1,98,000.00

Number of Machine Hours

            13,150

                                         15.06

Variable Machine Costs

2,51,840.00

Number of Machine Hours

            13,150

                                         19.15

Receiving department

                      2,44,500.00

Number of Receiving Orders

                     77

                                   3,175.32

Packing Department

                      1,06,500.00

Number of Packing Orders

                  145

                                      734.48

                      9,81,850.00

Calculation of Overheads allocated to the departments as per Activity Based Costing

Rose

Voilet

Costs

Number of Activities

Activity Rate $

Cost $

Number of Activities

Activity Rate $

Cost $

Variable Direct labour benefits $

                          33,550.00

                                               4.60

   1,54,330.00

                                   5,800.00

                  4.60

            26,680.00

Fixed Machine Costs

                          10,050.00

                                             15.06

   1,51,323.19

                                   3,100.00

                15.06

            46,676.81

Variable Machine Costs

                          10,050.00

                                             19.15

   1,92,470.87

                                   3,100.00

                19.15

            59,369.13

Receiving department

                                  53.00

                                       3,175.32

   1,68,292.21

                                         24.00

          3,175.32

            76,207.79

Packing Department

                                  97.00

                                           734.48

      71,244.83

                                         48.00

             734.48

            35,255.17

Total Overhead Cost $

7,37,661.10

2,44,188.90

Income Statement as per Activity Based Costing

Rose

Voilet

Total $ (54,500 Units)

Per Unit $

Total $ (10,900 Units)

Per Unit $

Sales

                    53,41,000.00

                    98.00

                         9,04,700.00

                 83.00

Less: Variable Cost

Material

                    27,25,000.00

                    50.00

                         4,68,700.00

                 43.00

Labour

                      5,99,500.00

                    11.00

                             98,100.00

                   9.00

Direct Labour Benefits

                      1,54,330.00

                      2.83

                           26,680.00

                   2.45

Machine Costs

1,92,470.87

                      3.53

59,369.13

                   5.45

Total Variable Cost

                    36,71,300.87

                    67.36

                         6,22,664.24

                 59.89

Contribution = Sales - Total Variable Cost

                    16,69,699.13

                    30.64

                       16,69,699.13

                 23.11

Less: Fixed Cost (Overheads)

Machine Costs

                      1,51,323.19

                      2.78

                             46,676.81

                   4.28

Receiving department

                      1,68,292.21

                      3.09

                             76,207.79

                   6.99

Packing department

                          71,244.83

                      1.31

                             35,255.17

                   3.23

Total Fixed Cost

                      3,90,860.23

                      7.17

                         1,58,139.77

                 14.51

Profit = Contribution - Total Fixed Cost

                    12,78,838.90

                    23.46

                       12,78,838.90

                   8.60

Calculation of Break Even Point (in Units) as per Activity Based Costing

= Total Fixed Cost / Contribution per unit

For Rose =$3,90,860.23 / $30.64 = 12,757.92 packages

For Voilet =$1,58,139.77 / $23.11 = 6,844.22 packages